Ruth and Stve MFJ. Have wages of $86500. They have a 1099 DIV with $1000 of Ordinary Dividends in Box 1 and $1000 of Qualified dividends in Box 2. What please

does this cost in Tax.

A, $150 B. $200 C. $ 220 or D $ 300.

Thank you.
Answer #1

Answer:-

C) $220

Explanation:-

1000 × 22% = 220

The dividend tax rate you will pay on ordinary dividends is 22%. Qualified dividends, on the other hand, are taxed at the capital gains rates, which are lower

Note :-

Dividends can be classified either as ordinary or qualified. Whereas ordinary dividends are taxable as ordinary income, qualified dividends that meet certain requirements are taxed at lower capital gain rates. The payer of the dividend is required to correctly identify each type and amount of dividend for you when reporting them on your Form 1099-DIV for tax purposes.

On other hand qualified dividend = 1000×15% = 150.